n is an integer chosen at random from the set {2,5,6} and p another integer chosen at random from the set {6,9,10}. What is the probability that the two numbers n and p are even?
A) 2/9 B) 5/9 C) 4/9 D) 1/9 E) 7/9

Answer by stanbon(57203)

n is an integer chosen at random from the set {2,5,6} and p another integer chosen at random from the set {6,9,10}. What is the probability that the two numbers n and p are even?
--------------
These selections are independent events.
P(even and even) = P(even)*P(even) = (2/3)(2/3) = 4/9
=====================
Cheers,
Stan H.
=====================
A) 2/9 B) 5/9 C) 4/9 D) 1/9 E) 7/9